# The prescribed metric on the boundary of convex subsets of anti-de Sitter space with a quasi-circle as ideal boundary

### v1

@@ Abstract.
Let $h^{+}$ and $h^{-}$ be two complete, conformal metrics on the disc $\mathbb{D}$. Assume moreover that the derivatives of the conformal factors of the metrics $h^{+}$ and $h^{-}$ are bounded at any order with respect to the hyperbolic metric, and that the metrics have curvatures in the interval $\left(-\frac{1}{\epsilon},-1-\epsilon\right)$, for some $\epsilon>0$. Let $f$ be a quasi-symmetric map. We show the existence of a globally hyperbolic convex subset $\Omega$ (see Definition 3.1) of the three-dimensional anti-de Sitter space, such that $\Omega$ has $h^{+}$ (respectively $h^{-}$) as the induced metric on its future boundary (respectively on its past boundary) and has a gluing map $\Phi_{\Omega}$ (see Definition 4.5) equal to $f$.
Introduction
@@
The three-dimensional anti-de Sitter space $\mathbb{ADS}^{2,1}$ is the Lorentzian analogue of the hyperbolic space $\mathbb{H}^{3}$. That is, any three-dimensional Lorentzian manifold that has constant sectional curvature equal to $-1$ is locally modeled on $\mathbb{ADS}^{2,1}$. A manifold which is locally modeled on $\mathbb{ADS}^{2,1}$, and which is oriented and time-oriented, is called an $\mathbb{ADS}$ space-time. Particular $\mathbb{ADS}$ space-times, called globally hyperbolic anti-de Sitter manifolds, have attracted attention since the work of [Mes07] that showed their relation to Teichmüller theory and their similarities with quasi-Fuchsian manifolds.
@@
A maximal globally hyperbolic $\mathbb{ADS}$ 3-manifold $M$ is diffeomorphic to $S\times(0,1)$, where $S$ is a closed connected surface of genus bigger or equal to $2$. A maximal globally hyperbolic manifold $\mathbb{ADS}$ 3-manifold is always the quotient of the domain of dependence of a quasi-circle $C_{M}$ (see Definitions 2.1 and 2.1) by a representation of the form $(\rho_{1},\rho_{2}):\pi_{1}(S)\to\rm{PSL}(2,\mathbb{R})\times\rm{PSL}(2,\mathbb{R})$ where $\rho_{1}$ and $\rho_{2}$ are Fuchsian representations (see [Mes07] or [BS20, Section 5]). This gives a one-to-one correspondence between maximal globally hyperbolic $\mathbb{ADS}$ 3-manifolds diffeomorphic to $S\times(0,1)$ and $\mathcal{T}(S)\times\mathcal{T}(S)$, where $\mathcal{T}(S)$ is the Teichmüller space of $S$. We have a similar correspondence for quasi-Fuchsian manifolds, which is known as Bers’ Theorem (see [Ber60]). Also, a globally hyperbolic manifold $M$ contains a smallest non-empty geodesically convex subset $C(M)$ which is called the convex core of $M$. When $C(M)$ is not a totally geodesic two-dimensional manifold, it has a boundary $\partial C(M)$ that consists of the disjoint union of two hyperbolic surfaces pleated along a measured lamination (see [Mes07] or [BS20, Section 6]).
@@
It was proved in [Dia13] that given any two hyperbolic metrics on $S$, we can find a globally hyperbolic manifold $M$ that induces those metrics on the boundary of $\partial C(M)$. An analogous result is known for quasi-Fuchsian manifolds (it follows from the work of [EM87] and [Lab92]). However, in both globally hyperbolic manifolds and quasi-Fuchsian manifolds, we don’t know if the manifold that induces the given metrics on the boundary of the convex core is unique or not.
@@
Later, Tamburelli in [Tam18] proved that given any two metrics on $S$ that have curvatures strictly smaller than $-1$, we can find a globally hyperbolic $\mathbb{ADS}^{2,1}$ manifold $N$ which is diffeomorphic to $S\times[0,1]$ with smooth spacelike boundary (it is embedded in a maximal globally hyperbolic manifold) that induces the given two metrics on $S$ on the boundary of $N$. A similar result was proved by [Lab92] in the case of quasi-Fuchsian manifolds. Unlike the $\mathbb{ADS}^{2,1}$ case, in which we don’t know the uniqueness of the manifold $N$ that induces the given metrics on its boundary, Schlenker in [Sch06] showed the uniqueness in the hyperbolic case.
@@
The authors in [BDMS21] have also shown a partial analogue of the main theorem in [Tam18] (when we assume that the curvatures of the induced metrics on the boundary are constant), which prescribes the metrics on the smooth spacelike boundaries of a convex compact $\mathbb{ADS}^{2,1}$ globally hyperbolic manifold. In the next theorem we mean by $K$-surface a convex space-like surface (we mean by convex that the product of its principal curvatures is positive) that has constant Gaussian curvature equal to $K$.
@@
The main theorem of this paper extends the main theorem of [Tam18] to the setting given in [BDMS21]. Our theorem gives a positive answer to the first question of Section 1.7 of [Sch20] under the hypothesis that the derivatives of the conformal factors of the metrics $h^{\pm}$ are bounded with respect to the hyperbolic metric (we will give the precise definition later).
@@
We denote by $h_{-1}$ the conformal hyperbolic metric on the disc $\mathbb{D}$. Let $h=e^{2\rho}h_{-1}$ be a complete conformal metric on the disc $\mathbb{D}$.
We say that $h$ has bounded derivatives at order $p$ by $M_{p}>0$ on the disc $\mathbb{D}$ if any derivative of $\rho$ at order $p$ is bounded at any point of the disc $\mathbb{D}$ by $M_{p}$ with respect to the hyperbolic metric.
We say that $h$ has bounded derivatives if its derivatives are bounded at any order.
Note that any conformal metric on $\mathbb{D}$ that has constant negative Gaussian curvature satisfies this condition. Also, note that any metric which is invariant under a Fuchsian representation $\rho:\pi_{1}(S)\to\rm{PSL}(2,\mathbb{R})$, where $S$ is a closed hyperbolic surface, satisfies this boundedness condition. Now we state our main Theorem, for a definition of globally hyperbolic convex subset see Definition 3.1,
@@
Let $h^{+}$ and $h^{-}$ be two complete, conformal metrics on the disc $\mathbb{D}$ that have curvatures in an interval of the form $(-\frac{1}{\epsilon},-1-\epsilon)$, for some $\epsilon>0$. Assume moreover that there is a sequence $(M_{p})_{p\in\mathbb{N}}$ of positive real numbers such that any derivative of $h^{+}$ or $h^{-}$ of order $p$ is bounded by $M_{p}$ (bounded with respect to the hyperbolic metric, see Definition 0.3). Let $f$ be a normalized quasi-symmetric map. Then there exists a globally hyperbolic convex subset $\Omega\subset\mathbb{ADS}^{2,1}$ such that the induced metric on $\partial^{+}\Omega$ is isometric to $(\mathbb{D},h^{+})$, the induced metric on $\partial^{-}\Omega$ is isometric to $(\mathbb{D},h^{-})$, and the gluing map of $\Omega$ is equal to $f$.
@@ 1.1. Quasi-conformal maps
$f:X\to Y$ be a diffeomorphism. We define the Beltrami differential $\mu=\mu(f)$ by the equation $\frac{\partial f}{\partial\bar{z}}=\mu\frac{\partial f}{\partial z}$. We say that $f$ is $K$ quasi-conformal if the dilatation number $K(f)=\frac{1+\left|\mu\right|_{\infty}}{1-\left|\mu\right|_{\infty}}$ is less than or equal to $K$. Note that we don’t need $f$ to be a $C^{1}$ diffeomorphism to define the notion of quasi-conformal maps. In fact, all we need is for $f$ to be a homeomorphism between $X$ and $Y$ that has derivatives in the sense of distribution that are $L^{2}$. For more details see [LV73]. The following proposition is well known, see for example [Tel07]
@@
Any quasi-conformal homeomorphism $f:\mathbb{H}^{2}\to\mathbb{H}^{2}$ extends to a homeomorphism $\partial f:\partial_{\infty}\mathbb{H}^{2}\to\partial_{\infty}\mathbb{H}^{2}$
@@ 1.2. Quasi-isometries
Let $C\geq 0$. We say that $f$ is an $(A,B,C)$ quasi-isometry if it is a $(A,B)$ quasi-isometric embedding and it is $C$-dense, that is:
@@
If $(X,d_{X})$ and $(Y,d_{Y})$ are $\delta$-hyperbolic spaces, then any quasi-isometric embedding $f$ extends uniquely to a homeomorphism $\partial f:\partial_{\infty}X\to\partial_{\infty}Y$ of the visual boundary. The homeomorphism $\partial f$ is called a quasi-symmetric map and has many interesting properties. We will discuss the notion of quasi-symmetric maps next. It is worth to mention the following well known proposition, for a proof see for example [Tel07]
@@
Any quasi-conformal map $f:\mathbb{H}^{2}\to\mathbb{H}^{2}$ is a quasi-isometric embedding
@@ 1.3. Quasi-symmetric maps
For the secend implication we refer for example to [Hub16] or [Tel07].
@@ 1.4. Cross-ratio
A quadruple of points $(a,b,c,d)$ is called symmetric if $cr(a,b,c,d)=-1$, or equivalently if there is $g\in PSL(2,\mathbb{R})$ such that $(g(a),g(b),g(c),g(d))=(0,1,-1,\infty)$. The following characterisation of quasi-symmetric maps using cross-ratio is well known, see [Tel07] for example.
@@
Conversely for any $M\geq 1$ there exists $k\geq 1$ such that if $f:\partial_{\infty}\mathbb{H}^{2}\to\partial_{\infty}\mathbb{H}^{2}$ is an orientation preserving homeomorphism that satisfies
$-M\leq crf(a,b,c,d)\leq-\frac{1}{M}$
@@ 1.5. The universal Teichmüller space
$\mathcal{T}=\mathcal{QS}(\partial_{\infty}\mathbb{H}^{2})/PSL(2,\mathbb{R})$
@@
Suppose $Y$ is another Riemann structure on $S$, and let $f:X\to Y$ be a diffeomorphism. We can identify $X$ with $\mathbb{H}^{2}/\Gamma$ and $Y$ with $\mathbb{H}^{2}/\Gamma^{\prime}$, where $\Gamma$ and $\Gamma^{\prime}$ are Fuchsian groups. Let $\tilde{f}:\mathbb{H}^{2}\to\mathbb{H}^{2}$ be a lift of $f$. Since $f$ is compact, its lift $\tilde{f}$ is quasi-conformal. Hence, so is $\tilde{f}$, by Proposition 1.3. $\tilde{f}$ extends to a unique quasi-symmetric map $\partial\tilde{f}$. Note that adjusting $X$ or $Y$ by isotopy changes $\partial\tilde{f}$ up to Möbius translation. In conclusion, once we fix a base point $X$ in $\mathcal{T}(S)$, we obtain a unique (up to composition by Möbius maps) quasi-symmetric map for each $Y\in\mathcal{T}(S)$. This correspondence is an embedding because for each two Fuchsian representations $\rho_{1},\rho_{2}:\pi_{1}(S)\to PSL(2,\mathbb{R})$, there exists a unique quasi-symmetric map which is $\rho_{1}$-$\rho_{2}$ equivariant. As a consequence, each point $Y$ in $\mathcal{T}(S)$ corresponds to a unique normalized quasi-symmetric map.

### v2

# The Prescribed Metric on Convex Subsets of Anti-de Sitter Space with Quasi-Circle Ideal Boundaries

@@ Abstract.
Let $h^{+}$ and $h^{-}$ be two complete, conformal metrics on the disc $\mathbb{D}$. Assume moreover that the derivatives of the conformal factors of metrics $h^{+}$ and $h^{-}$ are bounded at any order with respect to the hyperbolic metric, and that the metrics have curvatures in the interval $\left(-\frac{1}{\epsilon},-1-\epsilon\right)$, for some $\epsilon>0$. Let $f$ be a quasi-symmetric map. We show the existence of a globally hyperbolic convex subset $\Omega$ (see Definition 4.1) of the three-dimensional anti-de Sitter space, such that $\Omega$ has $h^{+}$ (respectively $h^{-}$) as the induced metric on its future boundary (respectively on its past boundary) and has a gluing map $\Phi_{\Omega}$ (see Definition 5.7) equal to $f$.
1. Introduction
@@
The three-dimensional anti-de Sitter space $\mathbb{ADS}^{2,1}$ is the Lorentzian analogue of the hyperbolic space $\mathbb{H}^{3}$. That is, any three-dimensional Lorentzian manifold that has constant sectional curvature equal to 1-1 is locally modeled on $\mathbb{ADS}^{2,1}$. A manifold which is locally modeled on $\mathbb{ADS}^{2,1}$, and which is oriented and time-oriented, is called an $\mathbb{ADS}$ space-time. Particular $\mathbb{ADS}$ space-times, called globally hyperbolic anti-de Sitter manifolds, have attracted attention since the work of [Mes07] that showed their relation to Teichmüller theory and their similarities with quasi-Fuchsian manifolds.
@@
A maximal globally hyperbolic $\mathbb{ADS}$ 3-manifold $M$ is diffeomorphic to $S\times(0,1)$, where $S$ is a closed connected surface of genus greater or equal to $2$. A maximal globally hyperbolic manifold $\mathbb{ADS}$ 3-manifold is always the quotient of the domain of dependence of a quasi-circle $C_{M}$ (see Definitions 3.1 and 3.1) by a representation of the form $(\rho_{1},\rho_{2}):\pi_{1}(S)\to\rm{PSL}(2,\mathbb{R})\times\rm{PSL}(2,\mathbb{R})$ where $\rho_{1}$ and $\rho_{2}$ are Fuchsian representations (see [Mes07] or [BS20, Section 5]). This gives a one-to-one correspondence between maximal globally hyperbolic $\mathbb{ADS}$ 3-manifolds diffeomorphic to $S\times(0,1)$ and $\mathcal{T}(S)\times\mathcal{T}(S)$, where $\mathcal{T}(S)$ is the Teichmüller space of $S$. We have a similar correspondence for quasi-Fuchsian manifolds, which is known as Bers’ Theorem (see [Ber60]). Also, a globally hyperbolic manifold $M$ contains a smallest non-empty geodesically convex subset $C(M)$ which is called the convex core of $M$. When $C(M)$ is not a totally geodesic two-dimensional manifold, it has a boundary $\partial C(M)$ which consists of the disjoint union of two hyperbolic surfaces pleated along a measured lamination (see [Mes07] or [BS20, Section 6]).
@@
It was proved in [Dia13] that given any two hyperbolic metrics on $S$, there exists a globally hyperbolic manifold $M$ that induces these metrics on the boundary $\partial C(M)$. An analogous result is known for quasi-Fuchsian manifolds (it follows from the works of [EM87] and [Lab92]). However, in both the globally hyperbolic and quasi-Fuchsian settings, It is not known whether the manifold that induces the given metrics on the boundary of the convex core is unique. Nevertheless, Prosanov proved the uniqueness of the manifold for an open and dense subset of the deformation space of quasi-Fuchsian manifolds (see [Pro22]).
For quasi-Fuchsian manifolds, Thurston conjectured that there is a one-to-one correspondence between the space of quasi-Fuchsian manifolds and the bending laminations of their convex core boundaries. That conjecture has been proved. Indeed, Bonahon and Otal (see [BO05]) showed that any two measured laminations on $S$ that fill the surface and do not have a closed leaf with weight greater than or equal to $\pi$ can be realized as the bending laminations of the convex core boundary of some quasi-Fuchsian manifold, which is homeomorphic to $S\times(0,1)$. Later, Dular and Schlenker (see [DS24b]) showed that if a pair of measured laminations on $S$ arises as the bending laminations of the boundary of the convex core of a quasi-Fuchsian manifold $Q$, then $Q$ is uniquely determined. In other words, the bending laminations uniquely determine the quasi-Fuchsian manifold.
For globally hyperbolic $\mathbb{ADS}^{2,1}$ manifolds, a similar result to the one given by Bonahon and Otal was obtained by Bonsante and Schlenker (see [BS12]). However, it is still unknown whether the bending laminations on the boundary of the convex core uniquely determine the globally hyperbolic manifold, exept if the bending lamination is small enough in a sense given in [BS12].
@@
Later, Tamburelli in [Tam18] proved that given any two metrics on $S$ that have curvatures stricly smaller than 1-1, we can find a globally hyperbolic $\mathbb{ADS}^{2,1}$ manifold $N$ which is diffeomorphic to $S\times[0,1]$ with smooth space-like boundary (it is embedded in a maximal globally hyperbolic manifold, see Figure 1) that induces the given two metrics on $S$ on the boundary of $N$. A similar result was proved by [Lab92] in the case of quasi-Fuchsian manifolds. Unlike the $\mathbb{ADS}^{2,1}$ case, in which we don’t know the uniqueness of the manifold $N$ that induces the given metrics on its boundary, Schlenker in [Sch06] showed the uniqueness in the hyperbolic case.
Figure 1. In other words, Tamburelli has shown that for any closed surface $S$ with genus greater than or equal to $2$, and for any Riemannian metrics $h^{+}$ and $h^{-}$ on $S$ with curvatures strictly smaller than 1-1, there exists a maximal globally hyperbolic manifold $M$ homeomorphic to $S\times(0,1)$ and a submanifold NN of $M$ that has the same homotopy type as $S$ and is diffeomorphic to $S\times[0,1]$, such that the boundary of NN consists of two disjoint space-like surfaces on which $M$ induces a metric on $S\times\left\{0\right\}$ homotopic to $h^{+}$ and a metric on $S\times\left\{1\right\}$ homotopic to $h^{-}$.
@@
The authors in [BDMS21] have also shown a partial analogue of the main theorem in [Tam18] (when we assume that the curvatures of the induced metrics on the boundary are constant), which prescribes the metrics on the smooth space-like boundaries of a convex compact $\mathbb{ADS}^{2,1}$ globally hyperbolic manifold. In the next theorem we mean by $K$-surface a convex space-like surface (we mean by convex that the product of its principal curvatures is positive) that has constant sectional curvature equal to $K$.
@@
The main theorem of this paper extends the main theorem of [Tam18] to the setting given in [BDMS21]. Our theorem gives a positive answer to the first question of Section 1.7 of [Sch20] under the hypothesis that the derivatives of the conformal factors of the metrics $h^{\pm}$ are bounded with respect to the hyperbolic metric.
@@
We denote by $h_{-1}$ the hyperbolic metric on the unit disc $\mathbb{D}$, written in conformal form as $h_{-1}=\frac{4\,|dz|^{2}}{(1-|z|^{2})^{2}}$. Let $h=e^{2\rho}h_{-1}$ be a complete conformal metric on $\mathbb{D}$, where $\rho:\mathbb{D}\to\mathbb{R}$ is a smooth function.
We say that $h$ has bounded derivatives of order $p$ if there exists a constant $M_{p}>0$ such that all derivatives of $\rho$ of order $p$, when measured with respect to the hyperbolic metric $h_{-1}$, are uniformly bounded on the disc by $M_{p}$, independently of the point $z\in\mathbb{D}$.
In terms of partial derivatives, this means that for any $p\in\mathbb{N}$ and for every multi-index $\alpha=(\alpha_{1},\alpha_{2})\in\mathbb{N}^{2}$ with $|\alpha|=p$, we have :
$\left|D^{\alpha}\rho(z)\right|\leq M_{p}(1-|z|^{2})^{p}\quad\text{for all }z\in\mathbb{D},$
where $D^{\alpha}=\partial_{x}^{\alpha_{1}}\partial_{y}^{\alpha_{2}}$ and $z=x+iy$. The factor $(1-|z|^{2})^{p}$ reflects that the derivatives are measured using the norm induced by the hyperbolic metric.
We say that $h$ has bounded derivatives if this property holds for every order $p\in\mathbb{N}$.
Note that any conformal metric on $\mathbb{D}$ that has constant negative sectional curvature satisfies this condition. Also, note that any metric which is invariant under a Fuchsian representation $\rho:\pi_{1}(S)\to\rm{PSL}(2,\mathbb{R})$, where $S$ is a closed hyperbolic surface, satisfies this boundedness condition. Now we state our main Theorem, for a definition of globally hyperbolic convex subset see Definition 4.1 (see also Figure 2).
@@
Let $h^{+}$ and $h^{-}$ be two complete, conformal metrics on the disc $\mathbb{D}$ that have curvatures in an interval of the form $(-\frac{1}{\epsilon},-1-\epsilon)$, for some $\epsilon>0$. Assume moreover that there is a sequence $(M_{p})_{p\in\mathbb{N}}$ of positive real numbers such that any derivative of $h^{+}$ or $h^{-}$ of order $p$ is bounded by $M_{p}$. Let $f$ be a normalized quasi-symmetric map. Then there exists a globally hyperbolic convex subset $\Omega\subset\mathbb{ADS}^{2,1}$ such that the induced metric on $\partial^{+}\Omega$ is isometric to $(\mathbb{D},h^{+})$, the induced metric on $\partial^{-}\Omega$ is isometric to $(\mathbb{D},h^{-})$, and the gluing map of $\Omega$ is equal to $f$.
@@
Figure 2. The connected components of $\partial\Omega\cap\mathbb{ADS}^{2,1}$, denoted by $\partial^{+}\Omega$ and $\partial^{-}\Omega$, are isometric to $(\mathbb{D},h^{+})$ and $(\mathbb{D},h^{-})$, respectively. The isometries $V^{+}$ and $V^{-}$ can be extended to the ideal boundary and then define the map $(\partial V^{-})^{-1}\circ\partial V^{+}$ (see Section 5), which is referred to as the gluing map of $\Omega$. In Theorem 1.4 this gluing map is equal to $f$.
@@ 2.1. Quasi-conformal maps
$f:X\to Y$ be an orientation preserving diffeomorphism. We define the Beltrami differential $\mu=\mu(f)$ by the equation $\frac{\partial f}{\partial\bar{z}}=\mu\frac{\partial f}{\partial z}$. We say that $f$ is $K$ quasi-conformal if the dilatation number $K(f)=\frac{1+\left|\mu\right|_{\infty}}{1-\left|\mu\right|_{\infty}}$ is less than or equal to $K$. Note that we don’t need $f$ to be a $C^{1}$ diffeomorphism to define the notion of quasi-conformal maps. In fact, all we need is for $f$ to be a homeomorphism between $X$ and $Y$ that has derivatives in the sense of distribution that are $L^{2}$. For more details see [LV73]. The following proposition is well known, see for example [FM07].
@@
Any quasi-conformal homeomorphism $f:\mathbb{H}^{2}\to\mathbb{H}^{2}$ extends to a homeomorphism $\partial f:\partial_{\infty}\mathbb{H}^{2}\to\partial_{\infty}\mathbb{H}^{2}$.
@@ 2.2. Quasi-isometries
Let $C\geq 0$. We say that $f$ is an $(A,B,C)$ quasi-isometry if it is an $(A,B)$ quasi-isometric embedding and it is $C$-dense, that is:
@@
If $(X,d_{X})$ and $(Y,d_{Y})$ are $\delta$-hyperbolic spaces, then any quasi-isometric embedding $f$ extends uniquely to a homeomorphism $\partial f:\partial_{\infty}X\to\partial_{\infty}Y$ of the visual boundary. The homeomorphism $\partial f$ is called a quasi-symmetric map and has many interesting properties. We will discuss the notion of quasi-symmetric maps next. It is worth to mention the following well known proposition, for a proof see for example [FM07].
@@
Any quasi-conformal map $f:\mathbb{H}^{2}\to\mathbb{H}^{2}$ is a quasi-isometric embedding.
@@ 2.3. Quasi-symmetric maps
For the second implication we refer for example to [Hub16] or [FM07].
@@ 2.4. Cross-ratio
A quadruple of points $(a,b,c,d)$ is called symmetric if cr(a,b,c,d)=1cr(a,b,c,d)=-1, or equivalently if there is $g\in PSL(2,\mathbb{R})$ such that $(g(a),g(b),g(c),g(d))=(0,1,-1,\infty)$.
For an orientation preserving homeomorphism $f:\partial_{\infty}\mathbb{H}^{2}\to\partial_{\infty}\mathbb{H}^{2}$, we denote:
$crf(a,b,c,d):=cr(f(a),f(b),f(c),f(d)).$
The following characterisation of quasi-symmetric maps using cross-ratio is well known, see [FM07] for example.
@@
Conversely for any $M\geq 1$ there exists $k\geq 1$ such that if $f:\partial_{\infty}\mathbb{H}^{2}\to\partial_{\infty}\mathbb{H}^{2}$ is an orientation preserving homeomorphism that satisfies :
$M\leq crf(a,b,c,d)\leq-\frac{1}{M},$
@@ 2.5. The universal Teichmüller space
$\mathcal{T}=\mathcal{QS}(\partial_{\infty}\mathbb{H}^{2})/PSL(2,\mathbb{R}).$
@@
Suppose $Y$ is another Riemann structure on $S$, and let $f:X\to Y$ be a diffeomorphism. We can identify $X$ with $\mathbb{H}^{2}/\Gamma$ and $Y$ with $\mathbb{H}^{2}/\Gamma^{\prime}$, where $\Gamma$ and $\Gamma^{\prime}$ are Fuchsian groups. Let $\tilde{f}:\mathbb{H}^{2}\to\mathbb{H}^{2}$ be a lift of $f$. Since $S$ is compact, $f$ is quasi-conformal. Hence, so is $\tilde{f}$, by Proposition 2.3. $\tilde{f}$ extends to a unique quasi-symmetric map $\partial\tilde{f}$. Note that adjusting $X$ or $Y$ by isotopy changes $\partial\tilde{f}$ up to Möbius translation. In conclusion, once we fix a base point $X$ in $\mathcal{T}(S)$, we obtain a unique (up to composition by Möbius maps) quasi-symmetric map for each $Y\in\mathcal{T}(S)$. This correspondence is an embedding because for each two Fuchsian representations $\rho_{1},\rho_{2}:\pi_{1}(S)\to PSL(2,\mathbb{R})$, there exists a unique quasi-symmetric map which is $\rho_{1}$-$\rho_{2}$ equivariant (that is, for all $\gamma\in\pi_{1}(S)$ and all $x\in\partial_{\infty}\mathbb{H}^{2}$, we have $\rho_{1}(\gamma)\cdot f(x)=f(\rho_{2}(\gamma)\cdot x)$). As a consequence, each point $Y$ in $\mathcal{T}(S)$ corresponds to a unique normalized quasi-symmetric map.